### v5

@@ Abstract
We bridge quasi-experimental and structural approaches to develop a robust method for retrospective merger evaluation. First, we show that the standard difference-in-differences estimation equation can be interpreted as a “reduced form” of our structural equations, formally bridging these methodologies. Demand and cost parameters enable the construction of a robust measure of the merger price effect that remains identified even when the difference-in-differences approach is not, and can be decomposed into efficiency gains versus conduct changes among the remaining firms. Second, we enhance the structural approach by weighting each market’s contribution to the moment conditions using market- and time-specific synthetic weights when the number of treated markets is limited. Applying this methodology to three major airline mergers, we find that modest efficiency gains are entirely offset by increased coordination, particularly in later mergers. The synthetic GMM refinement sharpens these findings, uncovering anti-competitive effects that standard approaches miss.
@@ 1 Introduction
We apply our method to evaluate three major airline mergers: Delta-Northwest (2008), United-Continental (2010), and American-US Airways (2013). Airline industry has been widely studied in academia (Borenstein, 1990; Kim and Singal, 1993; Peters, 2006; Olley and Town, 2018; Carlton et al., 2019; Das, 2019; Li et al., 2022; Eizenberg and Zvuluni, 2024; Olsen et al., 2024; Remer and Orchinik, 2024; Bruegge et al., 2024) and legal proceedings (U.S. District Court, MA, 2023).
@@
Second, we develop our first bridge by showing that the DiD is the reduced form of structural demand and supply equations. This connection reveals why DiD may not identify merger effects because, unlike standard policy interventions, mergers involve strategic interactions among firms that correlate with unobserved (to researchers) factors that affect demand, induce competitor responses, and may even alter competitive conduct—each potentially invalidating DiD identification, consistent with Marx et al. (2024). Our structural framework addresses these limitations by explicitly modeling these economic forces and relying on instrumental variables for identification.
@@
We model firm behavior using a conduct parameter, following Bresnahan (1982), and allow it to change in response to a merger. If the conduct parameter increases, then the merger has a coordination effect. With appropriate instruments, such as cost shifters (potential competitor presence, operational characteristics) and demand shifters (network effects, market size), we can identify merger-induced efficiency and the change in conduct but not levels. Using our structural model, we can also construct a DiD-like but robust measure of the aggregate price effect of a merger, which has an intuitive economic interpretation: it equals efficiency gains adjusted by the cost pass-through rate.
@@
Our estimates reveal a concerning temporal progression. The DL-NW merger exhibits decreased coordination (0.141-0.141), but this change is statistically insignificant. However, subsequent mergers increase coordination, with AA-US demonstrating significant competitive deterioration as its conduct increases by $0.107$ (5% level). This escalating pattern validates the DOJ’s concerns about coordinated effects that led it to block the AA-US merger. These conduct changes largely offset modest efficiency gains, yielding neutral price effects. However, several of the estimates are imprecise because there are limited treated markets–a common challenge in merger evaluation, where only a handful of markets or units are typically treated.
@@
Third, we develop our second bridge to utilize the DiD approach in addressing such imprecision in structural estimates. In particular, we propose using market- and time-specific weights from synthetic DiD (Arkhangelsky et al., 2021) to weight each observation’s contribution to the moment conditions. These weights help make treated and control markets comparable (Abadie and Gardeazabal, 2003), which should improve the precision of the “synthetic GMM” estimates. [^3] While our synthetic GMM approach is motivated specifically by merger analysis with limited treatment markets, the general principle of reweighting observations to address sampling concerns has precedent. For instance, Manski and Lerman (1977); Manski and McFadden (1981); Imbens (1992) proposed weighting likelihood and moment contributions to correct for choice-based sampling bias in discrete choice models. Indeed, we find that after weighting, estimates have better precision. Moreover, they suggest anti-competitive impacts that standard approaches miss: UA-CO shows significant coordination increases (0.049, 10% level) with efficiency losses of \$13.5 per \$100 ticket, suggesting that the price increases price by \$12.9, representing a 5.7% increase relative to the sample mean of \$228 for treated markets before meger, while AA-US exhibits substantial coordination effects (0.100, 5% level) with no discernable price effects.
@@ 2 Data
We adopt the framework established by Carlton et al. (2019) to define our analytical parameters, including time, control, and treatment groups. Specifically, for each merger, the pre- and post-merger periods are defined, respectively, as the eight quarters preceding and following the merger approval. Specifically, for DL-NW, these are Q4 2006-Q3 2008 (pre) and Q1 2009-Q4 2010 (post); for UA-CO, these are Q3 2008-Q2 2010 (pre) and Q4 2010-Q3 2012 (post); and for AA-US, these are Q4 2011-Q3 2013 (pre) and Q1 2014-Q4 2015 (post).
Next, we discuss the selection of “treated” and “control” markets, an approach that the US court system accepts when evaluating merger cases. For each merger, we classify a city-pair market as a “treated” market if both merging firms provided nonstop service. [^4] A carrier is considered ‘nonstop’ in a city route if it conducts 10 nonstop operations (five round-trips).
To minimize any indirect effects of the merger from these selected markets, we exclude markets where either of the two merging parties had a substantial connecting presence. A carrier is deemed to have a substantial connecting presence in a market if it handles at least 10% of total passenger traffic through connections. Specifically, we remove all “connecting overlap” markets, where, prior to the merger, each party served at least 10% of passengers and jointly served at least 40% of passengers. We refer to the remaining markets as the treated markets.
Likewise, we classify a market as a “control” market if the merger does not affect its market structure. In other words, in such a market, only one carrier can be the merging party, so if it is a duopoly or a triopoly before the merger, it remains the same after the merger. Thus, the control markets have the same number of nonstop carriers as the overlap routes in the pre-merger period. As with the treated market, we also restrict our attention to control markets where, whenever applicable, the merging firm does not have a substantial connecting presence.
@@ 3 Quasi-Experimental Approach
We use an event-study approach to assess the impact of three airline mergers (Delta-Northwest, United-Continental, and American-US Airways) on fares, passenger volume, and seat availability using a difference-in-differences (DiD) strategy. The DiD strategy compares the difference in the average outcome before and after a merger for treated markets with the difference for control markets to get the average effect of a merger on the treated markets.
@@ 3.1 Empirical Specification
provided that observable controls, $X_{mt}$, evolve similarly across the two groups of markets. This condition requires that the expected change in unobservable factors be identical for both groups while allowing for different levels of unobservables between groups or periods.
@@
$\beta^{\text{DiD}}$ $=$ $\Big(\mathbb{E}[Y_{m,\text{post}}|\text{Treated}_{m}=1]-\mathbb{E}[Y_{m,\text{post}}|\text{Treated}_{m}=0]\Big)$
$-\Big(\mathbb{E}[Y_{m,\text{pre}}|\text{Treated}_{m}=1]-\mathbb{E}[Y_{m,\text{pre}}|\text{Treated}_{m}=0]\Big).$
@@ 3.2 DiD Estimates of Merger Effects
These findings suggest all three mergers improved consumer welfare through lower prices and expanded service. Our results align with Carlton et al. (2019), who used similar DiD fixed-effects specifications and concluded that these mergers generated consumer benefits.
@@ 3.3 Identification of DiD with Fixed Effects
While the DiD framework provides a useful approach for evaluating mergers, its identification is tenuous in merger contexts. To understand these challenges systematically, we can conceptualize the DiD “error” term in Equation (1) $\varepsilon_{mt}$ as reflecting underlying economic forces that drive market outcomes. As it will become clearer later when we study the structural approach in Section 4, this error term can be decomposed into demand-side and supply-side components, making the connection between DiD identification failures and economic fundamentals transparent.
Endogenous Selection. The first concern arises from the strategic nature of merger decisions. The parallel trends assumption essentially requires that, conditional on observables ($X_{mt}$) and fixed effects ($\delta_{m}$, $\delta_{t}$), the selection of treated markets is “as good as random” with respect to future outcome trends. However, airlines evaluate potential merger partners based on network characteristics and anticipated market trajectories, creating two types of selection. Selection on demand occurs when merger decisions correlate with demand trends—airlines might merge in declining markets or, conversely, in growing markets to capitalize on expansion opportunities. Selection on supply occurs when merger decisions correlate with cost trajectories—airlines might target markets where joint operations would reduce costs or where standalone operations face rising expenses. Either form of selection invalidates the parallel trends assumption.
@@
Strict Exogeneity. An additional concern arises from the temporal structure of merger analysis. The fixed effects DiD estimator requires strict exogeneity: $\mathbb{E}[\varepsilon_{mt}\times I_{ms}]=0$ for all $m,t,s$. This assumption is particularly fraught when markets exhibit differential time trends. Consider the case where $\varepsilon_{mt}=\alpha_{m}+\kappa_{m}\cdot t+\nu_{mt}$, with $\nu_{mt}$ representing idiosyncratic noise and $\kappa_{m}$ capturing market-specific linear trends. If treated and control markets have different trend parameters, i.e., $\mathbb{E}[\kappa_{m}|\mathrm{Treated}_{m}=1]\neq\mathbb{E}[\kappa_{m}|\mathrm{Treated}_{m}=0]$, then bias grows linearly with time.
Specifically, when using a four-quarter pre- and post-merger window, the average bias across post-merger quarters equals $2.5(\kappa_{1}-\kappa_{0})$, where $\kappa_{1}$ and $\kappa_{0}$ represent average trend parameters for treated and control markets, respectively. Therefore, even a slight deviation from the pre-existing trend can result in substantial errors.
@@ 3.4 DiD with Market Trends
We implement two methods. Method 1 uses the entire sample period to estimate market-specific trends. Method 2 uses a two-step procedure that first estimates trends using only pre-merger data, and then extrapolates these trends to estimate the merger effects on detrended outcomes. For each method, we consider two variants: differential trends between treated and control groups (Panel A) and market-specific trends (Panel B). Table 3 shows the results.
@@
The previously positive passenger effects either disappear or reverse. Individual merger results show mixed patterns: DL-NW consistently shows significant passenger decreases, while AA-US shows increases in some specifications. However, the positive effect on seat capacity remains robust across all specifications, though magnitudes vary considerably. In comparison, baseline estimates indicated a 16.53% to 29.43% capacity increase, and trend-adjusted estimates (Method 1) range from 6.39% to 16.88%. Method 2 consistently produces larger capacity effects than Method 1.
The persistence of capacity expansion, coupled with the reversal in price and passenger effects, suggests changes in airline conduct rather than efficiency gains. The increase in offered seats, combined with insignificant or negative passenger effects, implies substantial reductions in load factors. This pattern may indicate strategic capacity deployment to maintain market presence rather than demand-driven expansion. The positive price effects in many trend-adjusted specifications are consistent with this interpretation.
@@ 3.5 DiD Estimates using First Differences
First differencing offers several advantages over the previous approaches. We find that outcome variables are non-stationary in levels but become stationary (Im et al., 2003) after first-differencing, addressing potential spurious regression concerns. More importantly, first differences require only that period-to-period changes in trends be similar across treated and control markets, accommodating arbitrary nonlinear trend differences between groups.
However, using first differences to estimate merger effects introduces a new challenge. Standard DiD estimates of (1) using the first-difference method capture only immediate post-merger effects, as the first-difference of the treatment indicator $\Delta I_{mt}$ equals one only in the period immediately following the merger. This timing limitation is particularly problematic for mergers, where the post-merger integration of operations and network optimization may not be immediate. Consequently, the realization of cost efficiencies may unfold gradually over multiple quarters or years following the merger. This pattern has been documented across several industries (Focarelli and Panetta, 2003; Gugler and Siebert, 2007; Ashenfelter and Hosken, 2010; Ashenfelter et al., 2013) and is likely applicable to airline mergers as well.
To address this limitation while maintaining the advantages of first differences, we allow for the gradual realization of merger effects over time. We estimate:
@@
where $I_{m,t,k}=\text{Post-Merger}_{t,k}\times\text{Treated}_{m}$ and $\text{Post-Merger}_{t,k}$ equals one if by period $t$ it has been at least $k$ periods since the merger. This modelling approach allows each $\beta^{\texttt{DiD}}_{k}$ to vary over time, enabling us to trace how merger effects evolve as integration progresses.
@@
The parallel trends assumption for first differences requires that counterfactual trends be identical across treated and control markets: $\mathbb{E}[\Delta\varepsilon_{mt}|\text{Treated}_{m}=1]=\mathbb{E}[\Delta\varepsilon_{mt}|\text{Treated}_{m}=0]$. While fixed effects require that the change in trend levels be the same, first differences require that period-to-period changes in trends be the same, which is substantially weaker. This approach accommodates arbitrary nonlinear trends, including different trend levels, growth rates, or functional forms, as long as the period-to-period changes are similar across groups.
@@
We extend our analysis to $T=8$ quarters (two years) post-merger to balance the tradeoff between allowing sufficient time for efficiency gains to materialize and avoiding contamination from other market developments. This horizon maintains comparability with the fixed effects estimates in Table 2.
@@
Figure 1 presents the first-difference estimates of the merger effects with time-varying treatment effects, allowing us to track how merger impacts evolve as integration progresses. Each “Quarter k” captures the effect when a treated market transitions into the post-merger state in quarter $k$ following merger completion. This specification reveals how merger effects unfold over the course of twelve quarters following each merger.
We find little evidence of stable price reductions across any of the three mergers. The DL-NW and UA-CO mergers exhibit modest price impacts that remain largely neutral over the observation period, while the AA-US merger shows no clear downward trend. Thus any efficiency benefits from these mergers do not translate into meaningful price reductions.
@@
We assess this impression by examining average effects over different time horizons. Table 4 reveals that airline mergers generally fail to deliver the promised efficiency gains that would benefit consumers through lower prices during the critical two-year post-merger integration period. Across the first eight quarters, two of the three mergers (DL-NW and AA-US) exhibit neutral price effects, while only UA-CO shows a meaningful price reduction of 1.68%. The DL-NW merger demonstrates no significant price change (0.97% increase) over quarters 1–8, while the AA-US merger similarly shows neutral effects (0.53% increase).
The heterogeneous outcomes underscore that merger effects depend critically on market-specific factors and the integration strategies employed, challenging the efficiency rationale. Before examining the underlying mechanisms that drive these diverse outcomes, we summarize our findings thus far.
@@
Beyond these methodological concerns, quasi-experimental approaches face deeper conceptual limitations. DiD estimates a single average treatment effect, masking rich heterogeneity in merger impacts and providing no guidance about which mergers benefit consumers or under what conditions. We therefore turn to a structural approach to address these limitations and provide economic insights into merger mechanisms.

### v6

@@ Abstract
We bridge quasi-experimental and structural approaches for robust merger evaluation. First, we show that the difference-in-differences (DiD) equation is the “reduced form” of a structural model, where demand and cost parameters identify price effects of mergers even when the DiD approach faces identification challenges. Second, we propose a synthetic GMM approach by applying synthetic DiD weights to structural moment conditions to improve estimates when only a few treated markets are available. Applying this methodology to three airline mergers, we find modest efficiency gains entirely offset by increased coordination. The synthetic GMM refinement sharpens findings, uncovering anti-competitive effects standard approaches miss.
@@ 1 Introduction
We apply our method to evaluate three major airline mergers: Delta-Northwest (2008), United-Continental (2010), and American-US Airways (2013). Airline industry has been widely studied; e.g., Borenstein (1990); Kim and Singal (1993); Peters (2006); Olley and Town (2018); Carlton et al. (2019); Das (2019); Li et al. (2022); Eizenberg and Zvuluni (2024); Olsen et al. (2024); Bruegge et al. (2024); U.S. District Court, MA (2023).
@@
Second, we develop our first bridge by showing that the DiD is the reduced form of market-level demand and supply equations. This connection reveals why DiD may not identify merger effects because, unlike standard policy interventions, mergers involve strategic interactions among firms that correlate with unobserved (to researchers) factors that affect demand, induce competitor responses, and may even alter competitive conduct—each potentially invalidating DiD identification, consistent with Marx et al. (2024). [^3] Interestingly, Blundell and McCurdy (1999) also explored the structural assumptions underlying the DiD approach to study labor supply effects and relate the estimated DiD parameters to those needed for policy. Our structural framework addresses these limitations by explicitly modeling these economic forces and relying on instrumental variables for identification.
@@
We model firm behavior using a conduct parameter, following Bresnahan (1982), and allow it to change in response to a merger. If the conduct parameter increases, then the merger has a coordination effect. With appropriate instruments, such as cost shifters (potential competitor presence, operational characteristics) and demand shifters (network effects, market size), we can identify merger-induced efficiency and the change in conduct but not levels. Using our structural estimates, we can also construct a robust DiD-like measure of the price effect of a merger, which has an intuitive economic interpretation: it equals efficiency gains adjusted by the cost pass-through rate.
@@
Our estimates reveal a concerning temporal progression. The DL-NW merger exhibits decreased coordination (0.141-0.141), but this change is statistically insignificant. However, subsequent mergers increase coordination, with AA-US demonstrating significant competitive deterioration as its conduct increases by $0.107$ (at the 5% level). This escalating pattern validates the DOJ’s concerns about coordinated effects that led it to block the AA-US merger. These conduct changes offset modest efficiency gains, yielding neutral price effects. However, several of the estimates are imprecise because there are limited treated markets–a common challenge in merger evaluation, where only a handful of markets are typically “treated.”
@@
Third, we develop our second bridge to utilize the DiD approach in addressing such imprecision in structural estimates. In particular, we propose using market- and time-specific weights from synthetic DiD (Arkhangelsky et al., 2021) to weight each observation’s contribution to the moment conditions. These weights help make treated and control markets comparable (Abadie and Gardeazabal, 2003; Abadie and Vives-i-Bastida, 2022) and improve the precision of estimates. [^4] While our synthetic GMM approach is motivated specifically by merger analysis with limited treatment markets, the general principle of reweighting observations to address sampling concerns has precedent. For instance, Manski and Lerman (1977); Manski and McFadden (1981); Imbens (1992) proposed weighting likelihood and moment contributions to correct for choice-based sampling bias in discrete choice models. Indeed, we find that after weighting, estimates have better precision. Moreover, they suggest anti-competitive impacts that standard approaches miss: UA-CO shows significant coordination increases (0.049, 10% level) with efficiency losses of \$13.5 per \$100 ticket, suggesting that the price increases price by \$12.9, representing a 5.7% increase relative to the sample mean of \$228 for treated markets before meger, while AA-US exhibits substantial coordination effects (0.100, 5% level) with no discernable price effects.
@@ 2 Data
We adopt the framework of Carlton et al. (2019) to define our analytical parameters, including time, control, and treatment groups. Specifically, for each merger, the pre- and post-merger periods are defined as the eight quarters preceding and following the merger approval, respectively. Specifically, for DL-NW, these are Q4 2006-Q3 2008 (pre) and Q1 2009-Q4 2010 (post); for UA-CO, these are Q3 2008-Q2 2010 (pre) and Q4 2010-Q3 2012 (post); and for AA-US, these are Q4 2011-Q3 2013 (pre) and Q1 2014-Q4 2015 (post).
Next, we discuss the selection of the “treated” and “control” markets, an approach the US court system accepts in evaluating merger cases. For each merger, we classify a city-pair market as a “treated” market if both merging firms provided nonstop service. [^5] A carrier is considered ‘nonstop’ in a city route if it conducts 10 nonstop operations (five round-trips).
To minimize indirect effects of the merger in these selected markets, we exclude markets where either of the two merging parties had a substantial connecting presence. A carrier is deemed to have a substantial connecting presence in a market if it handles at least 10% of total passenger traffic through connections. Specifically, we remove all “connecting overlap” markets, where, prior to the merger, each party served at least 10% of passengers and jointly served at least 40% of passengers. We refer to the remaining markets that are affected by a merger as the treated markets.
Likewise, we classify a market as a “control” market if the merger does not affect its market structure. In other words, in such a market, at most only one carrier can be the merging party. So, if it is a duopoly or a triopoly before the merger, it remains the same afterward. Thus, the control markets have the same number of nonstop carriers as the overlap routes in the pre-merger period. As with the treated market, we also restrict our attention to control markets where, whenever applicable, the merging firm does not have a substantial connecting presence.
@@ 3 Quasi-Experimental Approach
We use an event-study approach to assess the impact of three airline mergers (Delta-Northwest, United-Continental, and American-US Airways) on fares, passenger volume, and seat availability using a difference-in-differences (DiD) strategy. The DiD strategy compares the difference in the average outcome before and after a merger for treated markets with the difference for control markets to estimate the average effect of the merger on the treated markets.
@@ 3.1 Empirical Specification
provided that observable controls, $X_{mt}$, evolve similarly across the two groups of markets. This condition requires that the expected change in unobservable factors be identical across groups, while allowing for different levels of unobservables between groups or periods.
@@
$\hskip-50.58878pt\beta^{\text{DiD}}=\Big(\mathbb{E}[Y_{m,\text{post}}|\text{Treated}_{m}=1]-\mathbb{E}[Y_{m,\text{post}}|\text{Treated}_{m}=0]\Big)-\Big(\mathbb{E}[Y_{m,\text{pre}}|\text{Treated}_{m}=1]-\mathbb{E}[Y_{m,\text{pre}}|\text{Treated}_{m}=0]\Big).$
@@ 3.2 DiD Estimates of Merger Effects
These findings suggest that all three mergers improved consumer welfare by lowering prices and expanding services. Our results align with Carlton et al. (2019), who used DiD fixed-effects specifications and concluded that these mergers generated consumer benefits.
@@ 3.3 Identification of DiD with Fixed Effects
While the DiD framework provides a useful approach for evaluating mergers, its identification is tenuous in merger contexts. To understand these challenges systematically, we can conceptualize the DiD “error” term in Equation (1) $\varepsilon_{mt}$ as reflecting underlying economic forces that drive market outcomes. As it will become clearer later (in Section 4) when we study the structural approach, this error term can be decomposed into demand-side and supply-side components, making the connection between DiD identification failures and economic fundamentals transparent.
Endogenous Selection. The first concern arises from the strategic nature of merger decisions. The parallel trends assumption essentially requires that, conditional on observables ($X_{mt}$) and fixed effects ($\delta_{m}$, $\delta_{t}$), the selection of treated markets is “as good as random” with respect to future outcome trends. However, airlines evaluate potential merger partners based on network characteristics and anticipated market trajectories, leading to two types of selections. Selection on demand occurs when merger decisions correlate with demand trends—airlines might merge in declining markets or, conversely, in growing markets to capitalize on expansion opportunities. Selection on supply occurs when merger decisions correlate with cost trajectories—airlines might target markets where joint operations would reduce costs or where standalone operations face rising expenses. Either form of selection invalidates the parallel trends assumption.
@@
Strict Exogeneity. An additional concern arises from the temporal structure of merger analysis. The fixed effects DiD estimator requires strict exogeneity: $\mathbb{E}[\varepsilon_{mt}\times I_{ms}]=0$ for all $m,t,s$. This assumption is particularly fraught when markets exhibit differential time trends. Consider the case where $\varepsilon_{mt}=\alpha_{m}+\kappa_{m}\cdot t+\nu_{mt}$, with $\nu_{mt}$ representing idiosyncratic noise and $\kappa_{m}$ capturing market-specific linear trends. If treated and control markets have different trend parameters, i.e., $\mathbb{E}[\kappa_{m}|\mathrm{Treated}_{m}=1]\neq\mathbb{E}[\kappa_{m}|\mathrm{Treated}_{m}=0]$, then bias grows linearly with time. Specifically, when using a four-quarter pre- and post-merger window, the average bias across post-merger quarters equals $2.5(\kappa_{1}-\kappa_{0})$, where $\kappa_{1}$ and $\kappa_{0}$ represent average trend parameters for treated and control markets, respectively. Therefore, even a slight deviation from the pre-existing trend can result in substantial errors.
@@ 3.4 DiD with Market Trends
We implement two methods. Method 1 uses the entire sample period to estimate market-specific trends. Method 2 uses a two-step procedure: first, it estimates trends using only pre-merger data, and then extrapolates these trends to estimate the merger effects on detrended outcomes. For each method, we consider two variants: differential trends between treated and control groups (Panel A) and market-specific trends (Panel B). Table 3 shows the results.
@@
The previously positive passenger effects either disappear or reverse. Individual merger results show mixed patterns: DL-NW consistently shows significant passenger decreases, while AA-US shows increases in some specifications. However, the positive effect on seat capacity remains robust across all specifications, though magnitudes vary considerably. In comparison, baseline estimates indicated a 16.53% to 29.43% increase in capacity, and trend-adjusted estimates (Method 1) range from 6.39% to 16.88%. Method 2 consistently produces larger capacity effects than Method 1.
The persistence of capacity expansion, coupled with reversals in prices and passenger traffic, suggests changes in airline conduct rather than efficiency gains. The increase in offered seats, combined with insignificant or negative passenger effects, implies substantial reductions in load factors. This pattern may indicate strategic capacity deployment to maintain market presence rather than demand-driven expansion. The positive price effects in many trend-adjusted specifications are consistent with this interpretation.
@@ 3.5 DiD Estimates using First Differences
First differencing offers several advantages over the previous approaches. We find that the outcome variables are non-stationary in levels but become stationary after first-differencing (Im et al., 2003), addressing potential spurious regression concerns. More importantly, first differences require only that period-to-period changes in trends be similar across treated and control markets, accommodating arbitrary nonlinear trend differences between groups.
However, using first differences to estimate merger effects introduces a new challenge. Standard DiD estimates of (1) using the first-difference method capture only immediate post-merger effects, as the first-difference of the treatment indicator $\Delta I_{mt}$ equals one only in the period immediately following the merger. This timing limitation is particularly problematic for mergers, where the post-merger integration of operations and network optimization may not be immediate. Consequently, the realization of cost efficiencies may unfold gradually over time following the merger. This pattern has been documented across several industries (Focarelli and Panetta, 2003; Gugler and Siebert, 2007; Ashenfelter and Hosken, 2010; Ashenfelter et al., 2013) and is likely applicable to airline mergers as well.
To capture the gradual realization of merger effects over time, we estimate:
@@
where $I_{m,t,k}=\text{Post-Merger}_{t,k}\times\text{Treated}_{m}$ and $\text{Post-Merger}_{t,k}$ equals one if by period $t$ it has been at least $k$ periods since the merger. This specification allows each $\beta^{\texttt{DiD}}_{k}$ to vary over time, enabling us to trace how merger effects evolve as integration progresses.
@@
The parallel trends assumption for first differences requires that counterfactual trends be identical across treated and control markets: $\mathbb{E}[\Delta\varepsilon_{mt}|\text{Treated}_{m}=1]=\mathbb{E}[\Delta\varepsilon_{mt}|\text{Treated}_{m}=0]$. While fixed effects require that the change in trend levels be the same, first differences require that period-to-period changes in trends be the same, which is a substantially weaker condition. This approach accommodates arbitrary nonlinear trends, including different trend levels, growth rates, or functional forms, as long as the period-to-period changes are similar across groups.
@@
We consider $T=8$ quarters (two years) post-merger to balance the tradeoff between allowing sufficient time for efficiency gains to materialize and avoiding contamination from other market developments. This horizon maintains comparability with the fixed effects estimates in Table 2.
@@
Figure 1 presents the first-difference estimates of the merger effects with time-varying treatment effects, allowing us to track how merger impacts evolve as integration progresses. Each “Quarter k” captures the effect when a treated market transitions into the post-merger state in quarter $k$ following merger completion. This specification shows how merger effects unfold over the eight quarters following each merger.
We find little evidence of stable price reductions across any of the three mergers. The DL-NW and UA-CO mergers exhibit modest price impacts that remain largely neutral over the observation period, while the AA-US merger shows no clear downward trend. Thus, any efficiency benefits from these mergers do not translate into meaningful price reductions.
@@
We assess this impression by examining average effects over eight quarters. Table 4 reveals that airline mergers generally fail to deliver the promised efficiency gains that would benefit consumers through lower prices during the critical two-year post-merger integration period. Across the first eight quarters, two of the three mergers (DL-NW and AA-US) exhibit neutral price effects, while only UA-CO shows a meaningful price reduction of 1.68%. The DL-NW merger shows no significant price change (0.97% increase) over quarters 1–8, while the AA-US merger shows similarly neutral effects (0.53% increase).
The heterogeneous outcomes underscore that merger effects depend critically on market-specific factors and the integration strategies employed. Before examining the mechanisms underlying these diverse outcomes, we summarize our findings thus far.
@@
Beyond these methodological concerns, quasi-experimental approaches face deeper conceptual limitations. DiD estimates a single average treatment effect, which masks the rich heterogeneity in merger impacts and provides no guidance on which mergers benefit consumers or under what conditions. We therefore turn to a structural approach to address these limitations and provide economic insights into merger mechanisms.